Consider the R2−R function f defined by f(x,y)= lnxy and the R−R2 function r defined by r(t) = (t2,et). Determine the value of (f◦r)′(1) by using the General Chain Rule


1
Expert's answer
2021-06-08T17:09:19-0400

"f(x,y)=\ln xy"

"r(t)=(t^2,e^t)"

"(f\space o\space r)=f(r(t))=\ln t^2e^t"

"(f\space o \space r)'=\dfrac{1}{t^2 e^t}\times(2te^t+e^tt^2)=\dfrac{2}{t}+1"

"(f\space o \space r)'(1)=2+1=3"